The R# Medical Centre undertook a survey to determine the average waiting time of patients in the surgery before they saw a doctor. From 12 responses, the average and standard deviation of the waiting times was found to be 14.1 and 11.84 minutes, respectively. Assuming the information was collected through a random sample and that the waiting time is approximately normally distributed, calculate a 98% confidence interval for the mean waiting time at the medical centre. State the lower bound correct to two decimal places.
Answer #1

soin: here for 12 responses D = 22 8 0 = 0.02 f average waiting time of Standasd deviation for response was found to be 14.18pl - talq2t Leo/2) = (1-4) 981. PL-tale L x-lo Ltd 12 ) =(1-2J987 510 P/x-stal Leol 7 +5 tal =(1-d98%. ... Ci-a) 981 for the